The probability of the movie you want being at the Kroger redbox is 67% and the probability that the BlockBuster machine has the movie you want is 76 %. The probability that both have the movie is 56%.
A. What is the probability that neither machine has the movie you want?
B. What is the probability that only Rebox has the movie you want?
Explain

Answer by ikleyn(52777)

.
The probability of the movie you want being at the Kroger redbox is 67% and the probability
that the BlockBuster machine has the movie you want is 76 %. The probability that both have the movie is 56%.
A. What is the probability that neither machine has the movie you want?
B. What is the probability that only Rebox has the movie you want?
~~~~~~~~~~~~~~~~~


(A)  I will solve it in two steps.


         P(at least one) = 67% + 76% - 56% = 87%.


         Hence  P(no one) = 1 - P(at least one) =  1 - 87% = 13%.      ANSWER



(B)  P(only Redbox) = 67% - 56% = 11%.      ANSWER

Solved.